# Autoregressive Video Generation without Vector Quantization

### v1

@@ 1 Introduction
Autoregressive large language models (LLMs) (Brown et al. (2020); Touvron et al. (2023)) have become a foundational architecture in natural language processing (NLP), exhibiting emerging capabilities in in-context learning and long-context reasoning. In autoregressive (AR) vision generation domain, prior approaches (Ramesh et al. (2021); Ding et al. (2021); Yu et al. (2022); Yan et al. (2021); Villegas et al. (2022); Kondratyuk et al. (2023); Wang et al. (2024)) typically transform images or video clips into a discrete-valued token space using vector quantization (Van Den Oord et al. (2017); Esser et al. (2021)), which are then flattened into sequences for token-by-token prediction. However, it is challenging for vector-quantized tokenizers to achieve high fidelity and high compression simultaneously. More tokens are required for high quality. Thus, the cost increases substantially with higher image resolutions or longer video sequences.
@@
In this work, we present NOVA, which addresses the issues above and enables autoregressive video generation with high efficiency. We propose to reformulate the video generation problem as a non-quantized autoregressive modeling of temporal frame-by-frame prediction and spatial set-by-set prediction. NOVA is inspired by Emu3 (Wang et al. (2024)) for autoregressive video and multimodal generation, and MAR (Li et al. (2024c)) for non-quantized autoregressive image generation, which utilizes non-quantized vectors as visual tokens and performs set-by-set autoregressive prediction. While both are non-quantized autoregressive approaches, it is non-trivial at all from MAR to NOVA: 1) NOVA solves the challenges including efficiency, scalability, and mask schedule when learning more complex text-to-image generation instead of class-to-image generation. 2) NOVA first predicts temporal frames sequentially and then predicts spatial sets within each frame. NOVA is the first to enable a non-quantized autoregressive model for video generation.
@@ 2.2 Autoregressive Models for Visual Generation
Raster-scan Autoregressive Models are typically implemented on the discrete-valued RGB pixels (Kalchbrenner et al. (2017); Reed et al. (2017)) or latent space (Esser et al. (2021); Van Den Oord et al. (2017)), analogous to their language counterparts (Radford et al. (2019); Anil et al. (2023)). Recent studies involve scalable autoregressive transformers to generate token sequences in the raster-scan order for image generation (Ramesh et al. (2021); Ding et al. (2021; 2022); Yu et al. (2022); Sun et al. (2024b)), and video generation (Yan et al. (2021); Kondratyuk et al. (2023); Nash et al. (2022)). Specifically, VAR (Tian et al. (2024)) introduces next-scale prediction to progressively process the token-by-token sequence across multiple resolutions, leading to improved image quality.
@@ 3.1 Rethinking autoregressive models for video generation
As mentioned above, we regard text-to-video generation and autoregressive (AR) model as the basic task and means, respectively. We briefly retrospect related technical background. There exist two types of AR video generation approaches: (1) Token-by-token generation via raster-scan order. These studies perform causal per-token prediction within video frame sequence (Kondratyuk et al. (2023)), and decode vision tokens sequentially following the raster scan ordering (Wang et al. (2024)), which is defined as follows:
@@ 4.1 Experiment Setup
Evaluation. We use T2I-CompBench (Huang et al. (2023)), GenEval (Ghosh et al. (2024)) and DPG-Bench (Hu et al. (2024) to assess the alignment between the generated images and text condition. We generate image samples for each of the original or rewritten (Wang et al. (2024)) text prompts. Each image sample has a resolution of 512$\times$512 or 1024$\times$1024. We use VBench (Huang et al. (2024)) to evaluate the capacity of text-to-video generation across 16 dimensions. For a given text prompt, we randomly generate 5 samples, each with a video size of 33$\times$768$\times$480. We employ classifier-free guidance (Ho & Salimans (2022)) with a value of 7.0 along with 128 autoregressive steps to enhance the quality of the generated images and videos in all evaluation experiments.
Table 2: Text-to-image evaluation on various benchmarks. The best and second-best results are in blue and green. The data is from Huang et al. (2023),Wang et al. (2024) and Esser et al. (2024b).
@@ 4.2 Main Results
NOVA outperforms existing text-to-image models with superior performance and efficiency. In Table 2, we compare NOVA with several recent text-to-image models, including PixArt-$\alpha$ (Chen et al. (2023)), SD v1/v2 (Rombach et al. (2022b), SDXL (Podell et al. (2023)), DALL-E2 (Ramesh et al. (2022)), DALL-E3 (Betker et al. (2023b)), SD3 (Esser et al. (2024b)), LlamaGen (Sun et al. (2024b) and Emu3 (Wang et al. (2024)). After text-to-image training, NOVA achieves state-of-the-art performance on the GenEval benchmark, especially in generating a specified number of targets. Notably, NOVA also achieves leading results on T2I-CompBench and DPG-Bench, excelling at both the small model scale and data scale (requiring only 16% training overhead of the best competitor PixArt-$\alpha$). Last but not least, our text-to-video model outperforms most specialized text-to-image models, e.g., SD v1/v2, SDXL and DALL-E2. This underscores the robustness and versatility of our model in multi-context scenarios, with text-to-video generation as the fundamental training task.
@@
NOVA rivals diffusion text-to-video models and significantly suppresses the AR counterpart. We emphasize that the current version of our NOVA is designed to generate videos at 33 frames and can extend video length through the pre-filling of recently generated frames. We perform a quantitative analysis comparing NOVA against open-source and proprietary text-to-video models. As shown in Table 3, despite its significantly smaller size (0.6B vs. 9B), NOVA remarkably outperforms CogVideo (Hong et al. (2022)) across a variety of text-to-video evaluation metrics. It also matches the latest SOTA model Emu3’s (Wang et al. (2024)) performance (80.12 vs. 80.96) with a significantly smaller size (0.6B vs. 8B). Additionally, we compared NOVA with state-of-the-art diffusion models. This includes both closed-source models such as Gen-2 (Runway (2023)), Kling (Kuaishou (2024)), and Gen-3 (Runway (2024)), as well as open-source alternatives like LaVie (Wang et al. (2023)), Show-1 (Zhang et al. (2023a)), AnimateDiff-v2 (Guo et al. (2024)), VideoCrafter-v2.0 (Chen et al. (2024a)), T2V-Turbo ( Li et al. (2024b)), OpenSora-v1.1 (Zheng et al. (2024)), OpenSoraPlan-v1.1/v1.2 (Lin et al. (2024)), and CogVideoX (Yang et al. (2024)). The results underscore the effectiveness of text-to-image pre-training within our generalized causal process. Notably, we have narrowed the gap between autoregressive and diffusion methods in modeling large-scale video-text pairs, enhancing both the quality and instruction-following capabilities of video generation. Moreover, NOVA demonstrates a substantial speed advantage over existing models in terms of inference latency.
@@ 4.3 Qualitative Results
High-fidelity image and high-fluency video. We present a qualitative comparison of current leading image generation methods in Figure 3. NOVA demonstrates strong visual quality and fidelity across a range of prompt styles, and excels in color attribute binding and spatial object relationships. We present text-to-video visualizations in Figure 4, which highlight NOVA’s ability to capture multi-view perspectives, smooth object motion, and stable scene transitions based on the provided text prompts.
Zero-shot generalization on video extrapolation. By pre-filling generated frames, NOVA can produce videos that surpass the training length. For example, by shifting both the text and BOV embeddings, we generate 5-second videos that are up to twice the original length, as shown in Figure 5. We observed that during video extrapolation, NOVA consistently preserves temporal consistency of subject across frames. For instance, when the prompt describes a dome top and a lantern room, the model accurately represents the lighting within the house and captures the transition of a sunset. This further underscores the advantages of causal modeling in long-context video generation tasks.
@@
Zero-shot generalization on multiple contexts. By pre-filling the reference image, NOVA can generate videos from images, either with or without accompanying text. In Figure 6, we provide a qualitative example. We show that NOVA can simulate realistic motions without text prompts. Moreover, when text is included, perspective movements appear more natural. This indicates that NOVA is able to capture the fundamental physics, such as interaction forces and fluid dynamics.
@@ 4.4 Ablation Study
Effectiveness of temporal autoregressive modeling. To highlight the advantages of temporal autoregressive modeling, we have facilitated spatial autoregressive to finish video generation task. Specifically, we modify the attention mask of the temporal layer to bidirectional attention, and randomly predict the entire video sequence using set by set prediction. We observe less subject movement in videos under the same training iterations (Figure 7). Additionally, in zero-shot generalization across various contexts or video extrapolation, the network output exhibited more artifacts and temporal inconsistencies. Furthermore, this approach is not compatible with kv-cache acceleration during inference, leading to a linear increase in latency with the number of video frames. This further demonstrates the superiority of causal modeling over multitask approaches for video generation.
@@ 5 Conclusion
Acknowledgement. We thank Fan Zhang for the valuable advice on Figure 2 of this work, and Xiaotong Li, Jinming Wu, Chengyue Wu, Zhen Li, Quan Sun, Yueze Wang, Jinsheng Wang and Xiaosong Zhang for the insightful discussions. We also acknowledge the support provided by other colleagues at the Beijing Academy of Artificial Intelligence (BAAI) towards this project.
@@ Appendix E Ablations on the impact of Temporal Autoregressive Modeling
Under the same settings, we evaluate VBench results in Table 5 with and without using TAM (Temporal Autoregressive Modeling) to highlight its significance. Our findings are summarized as follows: (1) Efficient Motion Modeling: We observed that the total score was marginally lower without TAM compared to NOVA (75.38 vs. 75.84), especially in the dynamic degree metric, which showed a more pronounced decline (11.38 vs. 23.27). We hypothesize that while bidirectional attention enhances model capacity, it requires more extensive data and longer training times to capture subtle motion changes compared to causal models. (2) Efficient Video Inference: Thanks to the kv-cache technology and frame-by-frame autoregressive processing, NOVA’s inference time is much faster compared to methods without TAM, with a greater speed advantage for longer videos.

### v2

@@ 1 Introduction
Autoregressive large language models (LLMs) (Brown et al. (2020); Touvron et al. (2023)) have become a foundational architecture in natural language processing (NLP), exhibiting emerging capabilities in in-context learning and long-context reasoning. In autoregressive (AR) vision generation domain, prior approaches (Ramesh et al. (2021); Ding et al. (2021); Yu et al. (2022); Yan et al. (2021); Villegas et al. (2022); Kondratyuk et al. (2023); Wang et al. (2024a)) typically transform images or video clips into a discrete-valued token space using vector quantization (Van Den Oord et al. (2017); Esser et al. (2021)), which are then flattened into sequences for token-by-token prediction. However, it is challenging for vector-quantized tokenizers to achieve high fidelity and high compression simultaneously. More tokens are required for a high quality. Thus, the cost increases substantially with higher image resolutions or longer video sequences.
@@
In this work, we present NOVA, which addresses the issues above and enables autoregressive video generation with high efficiency. We propose to reformulate the video generation problem as a non-quantized autoregressive modeling of temporal frame-by-frame prediction and spatial set-by-set prediction. NOVA is inspired by Emu3 (Wang et al. (2024a)) for autoregressive video and multimodal generation, and MAR (Li et al. (2024c)) for non-quantized autoregressive image generation, which utilizes non-quantized vectors as visual tokens and performs set-by-set autoregressive prediction. While both are non-quantized autoregressive approaches, it is non-trivial at all from MAR to NOVA: 1) NOVA solves the challenges including efficiency, scalability, and mask schedule when learning more complex text-to-image generation instead of class-to-image generation. 2) NOVA first predicts temporal frames sequentially and then predicts spatial sets within each frame. NOVA is the first to enable a non-quantized autoregressive model for video generation.
@@ 2.2 Autoregressive Models for Visual Generation
Raster-scan Autoregressive Models are typically implemented on the discrete-valued RGB pixels (Kalchbrenner et al. (2017); Reed et al. (2017)) or latent space (Esser et al. (2021); Van Den Oord et al. (2017)), analogous to their language counterparts (Radford et al. (2019); Anil et al. (2023)). Recent studies involve autoregressive transformers to generate token sequences in the raster-scan order for image generation (Ramesh et al. (2021); Ding et al. (2021; 2022); Yu et al. (2022); Sun et al. (2024b)), and video generation (Yan et al. (2021); Kondratyuk et al. (2023); Nash et al. (2022); Wang et al. (2024b)). Specifically, VAR (Tian et al. (2024)) introduces next-scale prediction to progressively process the token-by-token sequence across multiple resolutions, leading to improved image quality.
@@ 3.1 Rethinking autoregressive models for video generation
As mentioned above, we regard text-to-video generation and autoregressive (AR) model as the basic task and means, respectively. We briefly retrospect related technical background. There exist two types of AR video generation approaches: (1) Token-by-token generation via raster-scan order. These studies perform causal per-token prediction within video frame sequence (Kondratyuk et al. (2023)), and decode vision tokens sequentially following the raster scan ordering (Wang et al. (2024a)), which is defined as follows:
@@ 4.1 Experiment Setup
Evaluation. We use T2I-CompBench (Huang et al. (2023)), GenEval (Ghosh et al. (2024)) and DPG-Bench (Hu et al. (2024) to assess the alignment between the generated images and text condition. We generate image samples for each of the original or rewritten (Wang et al. (2024a)) text prompts. Each image sample has a resolution of 512$\times$512 or 1024$\times$1024. We use VBench (Huang et al. (2024)) to evaluate the capacity of text-to-video generation across 16 dimensions. For a given text prompt, we randomly generate 5 samples, each with a video size of 33$\times$768$\times$480. We employ classifier-free guidance (Ho & Salimans (2022)) with a value of 7.0 along with 128 autoregressive steps to enhance the quality of the generated images and videos in all evaluation experiments.
Table 2: Text-to-image evaluation on various benchmarks. The best and second-best results are in blue and green. The data is from Huang et al. (2023),Wang et al. (2024a) and Esser et al. (2024b).
@@ 4.2 Main Results
NOVA outperforms existing text-to-image models with superior performance and efficiency. In Table 2, we compare NOVA with several recent text-to-image models, including PixArt-$\alpha$ (Chen et al. (2023)), SD v1/v2 (Rombach et al. (2022b), SDXL (Podell et al. (2023)), DALL-E2 (Ramesh et al. (2022)), DALL-E3 (Betker et al. (2023b)), SD3 (Esser et al. (2024b)), LlamaGen (Sun et al. (2024b) and Emu3 (Wang et al. (2024a)). After text-to-image training, NOVA achieves state-of-the-art performance on the GenEval benchmark, especially in generating a specified number of targets. Notably, NOVA also achieves leading results on T2I-CompBench and DPG-Bench, excelling at both the small model scale and data scale (requiring only 16% training overhead of the best competitor PixArt-$\alpha$). Last but not least, our text-to-video model outperforms most specialized text-to-image models, e.g., SD v1/v2, SDXL and DALL-E2. This underscores the robustness and versatility of our model in multi-context scenarios, with text-to-video generation as the fundamental training task.
@@
NOVA rivals diffusion text-to-video models and significantly suppresses the AR counterpart. We emphasize that the current version of our NOVA is designed to generate videos at 33 frames and can extend video length through the pre-filling of recently generated frames. We perform a quantitative analysis comparing NOVA against open-source and proprietary text-to-video models. As shown in Table 3, despite its significantly smaller size (0.6B vs. 9B), NOVA remarkably outperforms CogVideo (Hong et al. (2022)) across a variety of text-to-video evaluation metrics. It also matches the latest SOTA model Emu3’s (Wang et al. (2024a)) performance (80.12 vs. 80.96) with a significantly smaller size (0.6B vs. 8B). Additionally, we compared NOVA with state-of-the-art diffusion models. This includes both the closed-source models (Runway (2023); Kuaishou (2024); Runway (2024)), as well as the open-source alternatives (Wang et al. (2023); Zhang et al. (2023a); Guo et al. (2024); Chen et al. (2024a); Li et al. (2024b); Zheng et al. (2024); Lin et al. (2024); Yang et al. (2024)).
@@ 4.3 Qualitative Results
High-fidelity image and high-fluency video. We present a qualitative comparison of current leading image generation methods in Figure 3. NOVA demonstrates strong visual quality and fidelity across a range of prompts. We present text-to-video visualizations in Figure 4, which highlight NOVA’s ability to capture multi-view perspectives, smooth object motion, and stable scene transitions.
Zero-shot generalization on video extrapolation. By pre-filling generated frames, NOVA can produce videos that surpass the training length. For example, by shifting both the text and BOV embeddings, we generate videos that are up to twice the original length, as shown in Figure 5.
@@
Zero-shot generalization on multiple contexts. By pre-filling the reference image, NOVA can generate videos from images, either with or without accompanying text. In Figure 6, we provide a qualitative example. We show that NOVA can simulate realistic motions without text prompts.
@@ 4.4 Ablation Study
Effectiveness of temporal autoregressive modeling. To highlight the advantages of temporal AR, we have facilitated spatial AR to finish video generation task. We observe less subject movement in videos under the same training iterations (Figure 7). Additionally, in zero-shot generalization across various contexts or video extrapolation, the network output exhibited more artifacts and temporal inconsistencies. Furthermore, this approach is not compatible with kv-cache acceleration during inference, leading to a linear increase in latency with the number of video frames.
@@ 5 Conclusion
Acknowledgement. We thank Fan Zhang for the valuable advice on Figure 2 of this work, and Xiaotong Li, Jinming Wu, Chengyue Wu, Zhen Li, Quan Sun, Yueze Wang, Jinsheng Wang and Xiaosong Zhang for the insightful discussions. We also acknowledge the support provided by other colleagues at the Beijing Academy of Artificial Intelligence (BAAI) towards this project. This work was supported by the National Key R&D Program of China (2022ZD0116302), in part by Hainan Provincial Natural Science Foundation of China under No.624LALH008, and in part by the Program for Youth Innovative Research Team of BUPT under No. 2023YQTD02.
@@ Appendix E Ablations on the impact of Temporal Autoregressive Modeling
Under the same experimental settings, we evaluated VBench results both with and without using TAM (Temporal Autoregressive Modeling) to highlight its significance. Our findings are summarized as follows: (1) Efficient Motion Modeling: We observed that the total score was marginally lower without TAM compared to NOVA (75.38 vs. 75.84), especially in the dynamic degree metric, which showed a more pronounced decline (11.38 vs. 23.27). We hypothesize that while bidirectional attention enhances model capacity, it requires more extensive data and longer training times to capture subtle motion changes compared to causal models. (2) Efficient Video Inference: Thanks to the kv-cache technology and frame-by-frame autoregressive processing, NOVA’s inference time is much faster compared to methods without TAM, with a greater speed advantage for longer videos.